# Evaluating LLMs’ Mathematical Reasoning in Financial Document Question Answering

### v2

@@ 1. Incorrect Extraction:
  • Missing/Incomplete Evidences: The model fails to extract all the necessary evidences which serve as ingredients to derive the final answer.
  • Wrong Evidences: The model extracts wrong values for variables as supporting evidences from the premise.
@@ 2. Incorrect Reasoning:
  • Deficit in Domain Knowledge: These errors occur when the model attempts to derive an answer to the posed question using a wrong formula for domain-specific measures, for eg. COGS, ROI etc.
  • Question Misinterpretation: These errors occur when the model interprets the question differently and provides responses that are not aligned with the intended query. Overall, the model’s outputs lack grounding in the original question posed to it.
@@ 3. Incorrect Calculation:
This variety of errors include those where the model commits mistakes due to calculation mistakes. This can be of two types.
  • Incorrect Instantiation: These include cases if the model extracts the right evidences, uses the right derivation formula but instantiates the variables incorrectly with the values resulting in an incorrect answer.
  • Precision Error: Language models employ mathematical algorithms for arithmetic operations, but their results may not always be perfectly accurate due to insufficient data pattern coverage or introduced biases during training. Consequently, they can sometimes generate outputs with slight inaccuracies or deviations from correct results. We show a detailed analysis in A.1.
The above categorization provides a nuanced understanding of the diverse challenges and shortcomings exhibited in different facets of mathematical reasoning. We observe that in a lot of cases, the error propagates because of a deficiency in domain knowledge. It is critical for both evidence extraction and reasoning. Despite possessing general domain knowledge owing to the massive amount of data these models have been pre-trained upon, these models may require explicit prompts to elicit the specific domain knowledge needed for a particular question. Furthermore, errors can arise due to the models’ limited proficiency in multi-step reasoning, especially in tackling complex questions involving multiple arithmetic operations in a sequence.
We give a quantitative measure of each type of errors for each of the 4 datasets we consider for our study in Table 3. We also provide examples corresponding to each error category in figures 10, 11, 12, 13, 14 and 15. Statistically, we find that reasoning errors contribute a significant chunk to the total number of errors. In case of complex hierarchical tables like that in Multihiertt, the model is found to struggle with extracting the right supporting evidences from the premise for a given question. Calculation errors can be taken care of if a third-party calculation tool is chained to the language model.
Dataset Error Type Perc. (%)
FinQA Extraction.C1 10.38
Extraction.C2 25
Reasoning.C1 25
Reasoning.C2 15.57
Calculation 24.06
ConvFinQA Extraction.C1 8.45
Extraction.C2 14.08
Reasoning.C1 19.72
Reasoning.C2 36.62
Calculation 21.13
TATQA Extraction.C1 13.79
Extraction.C2 31.03
Reasoning.C1 22.41
Reasoning.C2 5.17
Calculation 27.59
Multihiertt Extraction.C1 20.5
Extraction.C2 31.5
Reasoning.C1 15.5
Reasoning.C2 12
Calculation 20.5
Table 3: Error Analysis on Various Datasets. In this table, Extraction.C1: Missing Evidences, Extraction.C2: Wrong Evidences, Reasoning.C1: Insufficient Domain Knowledge, Reasoning.C2: Question Misinterpretation, Calculation: Instantiation and Precision errors
@@ 5 Analysis on Reasoning Annotations
1. Performance vs Number of Reasoning Steps
[…]
@@ 1. Performance vs Number of Reasoning Steps
Anomalies are observed in ConvFinQA, where accuracy improves with two reasoning steps, potentially due to questions referring to prior conversation turns. Anomalies like these warrant further investigation beyond this study’s scope.
2. Performance across Question Types
We explore the performance trend across distinct question categories, as defined in Table 1, to evaluate the models’ familiarity with various mathematical and financial concepts. Figure 3 demonstrates that EEDP exhibits a performance competitive to PoT for simpler datasets like FinQA, ConvFinQA and TATQA but has the potential to significantly improve the performance in case of a complex dataset like Multihiertt.
[…]
@@ 6 Conclusion
In conclusion, our study delved into LLMs’ mathematical reasoning in complex financial scenarios, assessing their adaptability to mixed structured tables and unstructured text. Through rigorous experimentation, we uncovered insights into their performance and limitations, presenting a tailored prompting technique that outperformed other baseline methods. Our findings advance understanding of LLMs’ abilities in tackling intricate mathematical tasks within semi-structured documents, suggesting directions for future research.
7 Limitations
@@
  1. 1.
    There are not many datasets dealing with numerical reasoning over semi-structured data apart from the ones from financial domain. As a future work, it would be interesting to similar analysis across various domains such as e-commerce, healthcare and scientific tables from research papers.
  2. 2.
    It is essential to acknowledge that there may be multiple factors influencing the performance of large language models while dealing with numerical reasoning over semi-structured data. In our analysis, we have focused on specific factors and trends, but it is important to recognize that the overall performance is affected by a multitude of variables. Marginalizing (observing the trend along one while keeping the rest as constants) or isolating a single factor is challenging and cannot be done with real-world data. Additionally, future investigations may benefit from simulating controlled scenarios on synthetic datasets to gain deeper insights into the impact of individual factors on model performance.
  3. 3.
    We emphasize our analysis on prominent models such as GPT-4, GPT-3.5-TURBO, and PaLM 2-540B due to their substantial size and capabilities. Notably, other open-sourced large language models like LLaMA 2-13B, MAmmoTH-13B and Mistral-7B-Instruct did not exhibit satisfactory performance in numerical reasoning over semi-structured data. This accentuates the need for exploring computationally feasible and cheaper models that can deliver remarkable performance in tasks involving numerical reasoning over heterogeneous source of information.
  4. 4.
    As a future work, it would be intriguing to integrate special position encodings (similar to those utilized in TAPAS) into state-of-the-art language models for every cell in the table. This integration would facilitate the extraction of information from the table, considering its structure more effectively.
  5. 5.
    As a future direction, models specifically tuned to answer to human queries over complex documents in multiple conversational turns (like that in ConvFinQA) should also be considered as it’s a challenge for language model’s to backtrack to their decisions which were made previously in the conversation.
8 Ethics Statement
@@ A.1 How proficient are LLMs in performing simple arithmetic operations?
Figure 4 shows that in case of relatively simpler arithmetic operations like addition and subtraction, the effect of order of magnitude of the operands is less profound as compared to that in case of harder arithmetic operations like multiplication and division. We observe the trend in the performance accuracy with the growing and diminishing orders of magnitude. We also observe bigger model such as GPT-4, GPT-3.5-TURBO and PaLM 2-540B perform much better on simpler addition, subtraction task in comparison to multiplication, division task.
@@
Figure 4: Accuracy of different arithmetic operations across different orders of magnitude
A.2 Dataset Coverage across Reasoning Annotations
@@
Figure 6: Sample distribution of Numerical & Tabular Reasoning Datasets partitioned by Question Concept Category types. Clockwise (from top-left): FinQA, ConvFinQA, Multihiertt, TATQA
Figure 8: Sample distribution of Multihiertt Dataset partitioned by depth of hierarchy of the gold evidence
Figure 9: Sample Distribution of Multihiertt partitioned by the percentage of empty cells in the supporting table
A.3 Categorization of LLM prompting errors in numerical tabular reasoning
Basel III Standardized Transitional Basel III Advanced Transitional (in millions) Dec 31, 2017 Dec 31, 2016 Dec 31, 2017 Dec 31, 2016 Regulatory capital CET1 capital \$184,375 \$179,319 \$184,375 \$179,319 Tier 1 capital(a) \$184,375 \$179,341 \$184,375 \$179,341 Total capital \$195,839 \$191,662 \$189,419 \$184,637 Assets Capital ratios(c) CET1 13.8% 13.7% (e) 15.0% Tier 1(a) 13.8% 13.7% (e) 15.0% Total 14.7% 14.6% (e) 15.4% Tier 1 leverage(d) 8.7% 8.6% 8.7% Question: What is the sum of CET1 capital, Tier 1 capital, and Total capital in 2017? (in million) Response #: Domain Knowledge: In banking, CET1 capital, Tier 1 capital, and Total capital are used to measure a bank’s financial health and stability. CET1 (Common Equity Tier 1) capital consists mainly of common shares and retained earnings. Tier 1 capital is the core capital mainly comprising shareholders’ equity and disclosed reserves. Total capital includes both Tier 1 capital and Tier 2 capital, which consists of hybrid capital instruments and subordinated debt. Gold Evidences: Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional CET1 capital on Dec 31, 2017 as \$184,375 million. Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional Tier 1 capital on Dec 31, 2017 as \$184,375 million. Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional Total capital on Dec 31, 2017 as \$195,839 million. Steps: 1. Add CET1 capital and Tier 1 capital. Response 1: \$184,375 + \$184,375 = \$368,750 2. Add #1 and Total capital. Response 2: #1 + \$195,839 = \$564,589 The final answer is \$564,589 million. Ground-Truth Steps: \$184,375 + \$184,375 + \$195,839 + \$184,375 + \$184,375 + \$189,419 = \$1,122,758 Reason: The model missed adding JPMorgan Chase Bank, N.A. Basel III Advanced Transitional CET1 capital (\$184,375), Tier 1 (\$184,375), and Total Capital (\$189,419) in 2017.
@@ A.3 Categorization of LLM prompting errors in numerical tabular reasoning
2013 2014 2015 2016 2017 Masco \$138.48 \$155.26 \$200.79 \$227.08 \$318.46 S&P 500 Index \$132.04 \$149.89 \$151.94 \$169.82 \$206.49 S&P Industrials Index \$140.18 \$153.73 \$149.83 \$177.65 \$214.55 S&P Consumer Durables & Apparel Index \$135.84 \$148.31 \$147.23 \$138.82 \$164.39 Question: What was the difference in percentage cumulative total shareholder return on Masco common stock versus the S&P 500 Index for the five-year period ended 2017? Response #: Domain Knowledge: Cumulative total shareholder return is a measure of the overall performance of a stock over a given period of time. It includes the change in the stock’s price as well as any dividends paid to shareholders during that time. Gold Evidences: The cumulative total shareholder return for Masco in 2017 is \$318.46. The cumulative total shareholder return for the S&P 500 Index in 2017 is \$206.49. Steps: 1. Divide the cumulative total shareholder return for Masco in 2017 (318.46) by the cumulative total shareholder return for the S&P 500 index in 2017 (206.49) Response 1: 318.46 / 206.49 = 1.5422 2. Subtract 1 from the quotient obtained in #1 Response 2: 1.5422 - 1 = 0.5422 3. Multiply the value obtained in #2 with 100 to find the percentage difference Response 3: 0.5422 * 100 = 54.22 The final answer is 54.22%. Ground-truth steps: [((318.46 - 100.0)/100.$0$ ((206.49 - 100.0)/100.0)*100.0] = 111.97 Reason: Wrong formula for return on investment

### v3

@@ 1. Incorrect Extraction:
  • Missing/Incomplete Evidences (E1): The model fails to extract all the necessary evidences which serve as ingredients to derive the final answer.
  • Wrong Evidences (E2): The model extracts wrong values for variables as supporting evidences from the premise.
@@ 2. Incorrect Reasoning:
  • Deficit in Domain Knowledge (R1): These errors occur when the model attempts to derive an answer to the posed question using a wrong formula for domain-specific measures, for eg. COGS, ROI etc.
  • Question Misinterpretation (R2): These errors occur when the model interprets the question differently and provides responses that are not aligned with the intended query. Overall, the model’s outputs lack grounding in the original question posed to it.
@@ 3. Incorrect Calculation:
This variety of errors include those where the model commits mistakes due to calculation mistakes. This can be of two types as described below.
  • Incorrect Instantiation (C1): These include cases if the model extracts the right evidences, uses the right derivation formula but instantiates the variables incorrectly with the values resulting in an incorrect answer.
  • Precision Error (C2): Language models employ mathematical algorithms for arithmetic operations, but their results may not always be perfectly accurate due to insufficient data pattern coverage or introduced biases during training. Consequently, they can sometimes generate outputs with slight inaccuracies or deviations from correct results. We show a detailed analysis in A.1.
Analysis: The above categorization provides a nuanced understanding of the diverse challenges and shortcomings exhibited in different facets of mathematical reasoning. We observe that in a lot of cases, the error propagates because of a deficiency in domain knowledge. It is critical for both evidence extraction and reasoning. Despite possessing general domain knowledge owing to the massive amount of data these models have been pre-trained upon, these models may require explicit prompts to elicit the specific domain knowledge needed for a particular question. Furthermore, errors can arise due to the models’ limited proficiency in multi-step reasoning, especially in tackling questions involving multiple arithmetic operations in a sequence.
We give a quantitative measure of each type of errors for each of the 4 datasets we consider for our study in Table 3. We also provide examples corresponding to each error category in figures 12, 13, 14, 15, 16 and 17. Statistically, we find that reasoning errors contribute a significant chunk to the total number of errors. In case of complex hierarchical tables like that in Multihiertt, the model is found to struggle with extracting the right supporting evidences from the premise for a given question. Calculation errors can be taken care of if a third-party calculation tool (an external agent) is chained to the language model.
@@ 5 Analysis on Reasoning Annotations
1. Performance vs Number of Reasoning Steps.
@@
Anomalies are observed in ConvFinQA, where accuracy improves after greater than or equal to two reasoning steps, potentially due to questions referring to answers of prior conversation turns. Anomalies like these warrant further investigation beyond this study’s scope.
2. Performance across Question Types.
We analyze the performance trends across different question categories, as defined in Table 1, to assess the models’ understanding of various mathematical and financial concepts. Figure 3 shows that EEDP consistently performs better than or as well as PoT across all datasets. The improvement is particularly pronounced for PaLM-2-540B in all question categories.
3. Performance across Arithmetic Operations.
Figure 5 in Appendix A.1 shows that for relatively simpler arithmetic operations like addition and subtraction, the effect of order of magnitude of the operands is less profound as compared to harder arithmetic operations such as multiplication and division. We observe the trend in the performance accuracy with the growing and diminishing orders of magnitude. We also observe bigger and more capable models such as GPT-4, GPT-3.5-TURBO and PaLM 2-540B perform much better on simpler addition, subtraction task in comparison to multiplication, division task. For more details on refer to the Appendix A.1.
6 Other Related Works
6.1 LLMs on Mathematical Reasoning
6.2 Tabular Question Answering
Handling diverse input formats in question answering, including structured tables and visual data, poses challenges for language models. HybridQA Chen et al. (2020b) introduces questions requiring reasoning over tables and text. MultimodalQA Talmor et al. (2021) adds visual inference. Our focus is on multi-hop question answering over tables and text. TAPAS Herzig et al. (2020) tackles table-based questions without logical forms, while Tapex Liu et al. (2022) empowers generative models with table reasoning.
Models like FinQANet, TagOP, and MT2Net Chen et al. (2021b); Zhu et al. (2021); Zhao et al. (2022) employ a retriever module to extract supporting facts from input financial reports, followed by a reasoning module to derive the final answer. Retrieving relevant evidence and conducting reasoning both demand domain-specific knowledge, such as understanding financial transactions, identifying revenue trends, and interpreting complex investment statements. Thus, reliance on an external knowledge base becomes crucial for addressing the challenges of domain-specific multi-hop question answering.
6.3 Prompt Engineering
In-context Learning (ICL) equips Language Models (LLMs) with task examples and queries, enabling them to perform target tasks without updating model parameters (Brown et al., 2020; OpenAI, 2023). They excel in mathematical reasoning with few-shot prompts but struggle with more complex tasks. Methods like chain-of-thoughts (CoT) (Wei et al., 2022) have improved LLM performance by guiding them through intermediate reasoning steps. Enhancing multi-step reasoning involves two main approaches: improving in-context examples and obtaining better reasoning steps. Some focus on stable in-context example selection (Fu et al., 2023; Rubin et al., 2022; Lu et al., 2023). Others adopt a modular approach, using off-the-shelf tools (Schick et al., 2023), program of thoughts (PoT) (Chen et al., 2022a), or decomposition strategies (Zhou et al., 2023; Dua et al., 2022; Khot et al., 2023).
Our task requires complex multi-step reasoning across diverse information sources. LLMs, as demonstrated by (Chen, 2023), can reason over structured tables without explicit encoding. They also serve as versatile decomposers, breaking down extensive evidence and complex questions (Ye et al., 2023).
7 Key Takeaways
Our Contributions.
Our study aimed to delve into the mathematical reasoning abilities of LLMs within the context of financial documents where models are tasked with complex hybrid (table-text) question answering. This presents a significant challenge, requiring models not only to provide accurate numerical analysis but also to retrieve right supporting evidence tailored to specific question requirement. Moreover, it necessitates the extraction of necessary knowledge from the model’s pre-trained parameters to address queries.
Firstly, we meticulously annotate popular financial datasets, such as FinQA, ConvFinQA, TATQA, and Multihiertt, with detailed meta-information. This includes specifying mathematical operations, types of reasoning involved, table dimensions, question types, and the depth of table hierarchy. Furthermore, we conduct a manual error analysis to quantify error types across multiple LLMs. These detailed annotations are invaluable for analyzing various dimensions where LLM models encounter challenges. This, in turn, aids in the development of better prompting techniques such as EEDP, aimed at enhancing LLMs’ mathematical reasoning abilities. The resulting improvement in performance with EEDP across multiples datasets serves as compelling evidence of the effectiveness of this approach.
What did we learn?
Our analysis revealed that LLMs can accurately handle addition and subtraction tasks e.g. modle perform fairly when calculating total expenses or profits, but struggle with multiplication and division e.g. model performs poorly with questions requiring reasoning operations involving proportions, ratios, percentages, and division. Moreover, as the complexity of the data increases either through a higher absolute order scale or more decimal numbers, model performance degrades. Model performance also degrades with increasing number of reasoning steps and lengthy complex hierarchical table structures. e.g. in complex datasets with hierarchical structures such as Multiheirtt, TATQA, incorrect extraction leads to modeling errors. Similarly, on queries involving multiple conversational turns, such as those in ConvFinQA, model perform poorly due to reasoning failures, like misinterpreting multiple queries longer context. Across all models, incorrect reasoning and incorrect extraction consistently emerge as common sources of errors. For smaller models, even straightforward calculations, result in errors due to imprecise calculations.
EEDP vs other methods
(a.) EEDP vs PoT:
PoT enhances LLM inference with the use of variable names for the supporting values extracted from the premise and prompts the LLM to express their thought process in the form of programs. The model output is a program which is executed externally to derive the final answer. EEDP proposes to decompose a complex reasoning task into simple atomic steps whose solutions can be composed to give the final answer. In PoT, they don’t make the language model do the computation while in our case the language model not only outputs the reason but also computes the final answer. This distinction implies that PoT may have an inherent advantage over EEDP.
(b.) EEDP vs Decomposers:
The prompting strategy proposed by Ye et al. (2023) was originally designed for querying SQL tables, they use LLMs to break down evidence and questions for SQL interpreters. In contrast, our approach addresses more complex scenarios involving both tables and text, requiring advanced reasoning skills beyond the capability of standard SQL interpreters. Pruning a non-SQL table using this method can lead to significant information loss from the premise which can be a potential ingredient required to derive the final answer. Additionally, this is an expensive method as it requires 3X API calls as opposed to other prompting methods. Moreover, EEDP is a unified prompting strategy which integrates multiple solver elements into a single unified prompt for elicitation, extraction, decomposition and prediction.
@@ 8 Conclusion
In conclusion, our study delved into LLMs’ mathematical reasoning in complex financial scenarios, assessing their adaptability to mixed structured tables and unstructured text. Through rigorous experimentation, we uncovered insights into their performance and limitations, presenting a tailored prompting technique that outperformed other baseline methods. Our findings advance understanding of LLMs’ abilities in tackling intricate mathematical tasks within semi-structured documents, suggesting directions for future research. Please refer to appendix section A.3 for future work details.
Limitations
@@
Dataset Scarcity.
There are not many datasets dealing with numerical reasoning over semi-structured data apart from the ones from financial domain. As a future work, it would be interesting to similar analysis across various domains such as e-commerce, healthcare, sports and scientific tables from research papers, uncovering new challenges and insights. This expansion will enhance the applicability and impact of our research within the NLP community. However, creating tailored datasets for these domains presents a significant challenge.
For now to ensure a comprehensive evaluation of LLMs, we have integrated financial datasets that offer diverse challenges. For instance, Multihiertt evaluates model performance with intricate premise structures, providing insights into handling complex data hierarchies. ConvFinQA delves into the intricate chains of numerical reasoning within conversational question answering contexts, offering a unique perspective on dynamic data interpretation. Moreover, FinQA and TAT-QA encompass a wide array of reasoning types, with a significant portion requiring domain-specific knowledge, thereby broadening the evaluation spectrum.
Factors Isolation.
It is essential to acknowledge that there may be multiple factors influencing the performance of large language models while dealing with numerical reasoning over semi-structured data. In our analysis, we have focused on specific factors and trends, but it is important to recognize that the overall performance is affected by a multitude of variables. Marginalizing i.e. observing the trend along one while keeping the rest as constants or isolating a single factor is challenging and cannot be done with real-world data. Additionally, future investigations may benefit from simulating controlled scenarios on synthetic and counterfactual datasets to gain deeper insights into the impact of individual factors on model performance.
Modeling Improvement.
We emphasize our analysis on prominent models such as GPT-4, GPT-3.5-TURBO, and PaLM 2-540B due to their substantial size and capabilities. Notably, other open-sourced large language models like LLaMA 2-13B, MAmmoTH-13B and Mistral-7B-Instruct did not exhibit satisfactory performance in numerical reasoning over semi-structured data. For more detail about the the model choices refer to Appendix A.2. This accentuates the need for exploring computationally feasible and cheaper models that can deliver remarkable performance in tasks involving numerical reasoning over heterogeneous sources of information. Future experiments with ample computational resources may involve exploring larger open-source models like OLMo, Mixtral, and DBRX, which have been recently released.
Ethics Statement
@@ Acknowledgements
Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-20-1-0080. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. This work was partially funded by ONR Contract N00014-19-1-2620. Lastly, we extend our appreciation to the reviewing team for their insightful comments.
@@ A.1 How proficient are LLMs in performing simple arithmetic operations?
Analysis. Figure 5 illustrates that for simpler arithmetic operations like addition and subtraction, the impact of the order of magnitude of the operands is less significant compared to harder operations like multiplication and division. We observe a trend in performance accuracy with increasing and decreasing orders of magnitude. Larger models such as GPT-4, GPT-3.5-TURBO, and PaLM 2-540B perform significantly better on addition and subtraction tasks as compared to the multiplication and division tasks.
@@
Figure 5: Accuracy of different arithmetic operations across different orders of magnitude.
(a) Hierarchical Complexity
(b) Empty Cells
Figure 7: Sample distribution of Multihiertt Dataset partitioned by (a) hierarchical complexity of the gold evidence. (b) the percentage of empty cells in the supporting table.
A.2 Model Selection Criteria
Our model selection process was guided primarily by resource constraints and the timeframe of our research endeavor. We aimed to identify models that represented the state-of-the-art (SOTA), such as GPT-4, or those with a specific focus on mathematical reasoning, such as MAmmoTH, aligning with the parameters of our project. Here’s a breakdown:
  1. 1.
    Resource and Budget Constraints: Given our limited computing resources and budget, we meticulously selected models that could provide valuable insights within the confines of our project. Incorporating additional models would have been impractical given our resource limitations.
    The number of shots, indicated by #shots = k, was tailored to the context length of the model. Specifically, for models with a context length exceeding the input length, we standardized k to 4. For instance, we allocated 2 shots for models like LLaMA and MammoTH due to their constrained context length. However, for other models capable of accommodating larger contexts, we increased the number of shots to 4. Additionally, we used a temperature of 0 and topp = 1 for our experiments.
  2. 2.
    Models with Mathematical Capabilities: We prioritized models renowned for their advanced mathematical prowess, such as MAmmoTH, alongside state-of-the-art Large Language Models (LLMs) like GPT-4. Our goal was to gain deeper insights into the mathematical reasoning capabilities of cutting-edge models within the context of financial documents.
  3. 3.
    Better Prompting Approaches: Rather than focusing solely on model diversity, we concentrated on exploring a variety of prompting methods, particularly those aimed at enhancing mathematical reasoning. We believed this approach would yield more valuable insights into the performance of both LLMs and their associated prompting techniques in real-world financial tasks.
  4. 4.
    Excluding Underperforming Models: While we experimented with various models, such as Falcon-7B-Instruct and MPT-7B-Instruct, we found them to underperform significantly compared to models like LLaMA and Mistral. Consequently, we excluded them from detailed analysis. Future experiments with ample computational resources may involve exploring additional open-source models like OLMo, Mixtral, and DBRX, which have been recently released.
A.3 Other Modeling Techniques
Based on our research and the results obtained from our proposed method ’EEDP’, we do have several insights that could guide future model development:
  1. 1.
    Domain-Specific Pre-training: Our method "EEDP" suggests that LLMs could benefit from pre-training that focuses on extracting domain-specific knowledge. In the context of financial documents, for instance, this could involve training models on a corpus of financial texts, thereby enabling them to better understand and reason about financial concepts and terminology.
  2. 2.
    Knowledge Elicitation: The elicitation step in "EEDP" indicates the potential for designing LLMs that can elicit or extract relevant information from a given context more effectively. This could involve developing models that are better at identifying and focusing on key pieces of information in a document, which is crucial for accurately answering questions about the document.
  3. 3.
    Modular Modeling: Our research introduce a novel approach to the reasoning process, wherein it’s broken down into modular steps. In this methodology, Large Language Models (LLMs) handle different aspects of a task in distinct stages. This division potentially enhances the overall accuracy and efficiency of the model.
    For instance, the model might begin by eliciting domain-specific knowledge, then proceed to extract relevant information from the premise. Following this, it engages in reasoning about this information to answer a question and finally derives the answer, using the output from the preceding reasoning steps as a reference point.
    By potentially training individual expert models to handle each specific stage, we could optimize performance for each distinct task. This modular approach allows for specialized processing of each step, thereby improving the overall performance and interpretability of the final output.
  4. 4.
    Hierarchical Structure Understanding: Representing the input structure of the table in a better format to the LLM could be beneficial. One can also explore introducing special positional encodings, similar to those used in TAPAS, to serve as row and column IDs for each cell. This approach would differ from traditional positional encodings, which are designed to capture the inherently linear structure of textual data. TThis integration would facilitate the extraction of relvant information from the table correctly, considering its structure more effectively, avoiding information extraction errors. Another idea could be converting the premise containing the complex table and text into a common representation such as a knowledge graph. Furthermore, models specifically tuned to answer to human queries over complex documents in multiple conversational turns (like that in ConvFinQA) should also be considered, as it’s a challenge for language model’s to backtrack to their decisions that were made previously in the conversation.
A.4 Metadata Annotations Dataset Coverage
Figure 6 displays the dataset distribution for Multihiertt and FinQA based on table length (number of rows). Figure 7(a) (a) shows how we calculate hierarchical complexity for examples with multiple relevant rows at various hierarchical depths. For the distribution of missing information (empty cell proportions) across datasets, refer to Figure 7(a) (b). Figure 8 for the distribution of questions by reasoning steps. We define 12 mathematical concept categories, see Table 1) and annotate each question accordingly. The dataset coverage for these categories is shown in Figure 9. Our method for estimating hierarchy depth is shown in Figure 10. Figure 11 shows a example for EEDP strategy with one shot. Figure 11 shows a example for EEDP strategy with one shot.
@@ A.4 Metadata Annotations Dataset Coverage
Figure 9: Sample distribution of Numerical & Tabular Reasoning Datasets partitioned by Question Concept Category types. Clockwise (from top-left): FinQA, ConvFinQA, Multihiertt, TATQA.
Years Ended December 31, 2009 2008 (in millions, except percentages) Revenues       Management and financial advice fees \$1,234 \$1,339       Distribution fees \$1,733 \$1,912       Net investment income \$297 \$-43       Other revenues \$85 \$80       Total revenues \$3,349 \$3,288 Banking and deposit interest expense \$133 \$178 Total net revenues \$3,216 \$3,110 Expenses       Distribution expenses \$1,968 \$2,121       General and administrative expense \$1,282 \$1,138       Total expenses \$3,250 \$3,259 Pretax loss \$-34 \$-149 Question: What will Distribution fees reach in 2010 if it continues to grow at its current rate? (in millions)? Gold Evidences: Table shows Distribution fees of Years Ended December 31, 2009 (in millions, except percentages) is \$1,733. Table shows Distribution fees of Years Ended December 31, 2008 (in millions, except percentages) is \$1,912.
Hierarchical Complexity: max(4, 4) = 4
Figure 10: This illustration demonstrates how the "hierarchical complexity" is determined for each supporting piece of evidence. The hierarchical complexity corresponds to the number of levels or tiers of information structure within the table. The top level (1) encompasses the table itself, which contains information about the results of operations in a firm’s Advice & Wealth Management segment. The second level (2) includes columns specifying the years (2009, 2008), creating a substructure within the table. The third level (3) involves the column "(in millions, except percentages)," and the fourth level (4) encompasses the rows under categories like "Revenues," "Distribution fees," "Net investment income," and so on. To locate specific values like "\$1733" and "\$1912", a retriever module needs to navigate through these four levels.
Instruction: Please carefully analyze the provided information, perform necessary numerical calculations, and provide accurate answers to the given question using the provided data. Take into account the pre-text, table, and post-text when formulating your response. When performing numerical calculations, ensure you gather the required information and follow a step-by-step approach. 1. Elicit the most relevant domain fact or knowledge that might be useful for you to extract the right operands and forming the right approach to the problem. 2. Translate table data into sentences and identify gold evidence for answering the question. 3. Define a sequence of atomic operations (Add, Subtract, Divide, Multiply, Greater, Min, Max, Exp) which take into account only two operands at a time. Divide a complex task into a sequence of atomic operations defined above. 4. Finally, conclude as: The final answer is <answer>. Demonstration: “Effective Income Tax Rate”: A reconciliation of the United States federal statutory income tax rate to our effective income tax rate is as follows: In 2019 and 2018 we had pre-tax losses of \$19,573 and \$25,403, respectively, which are available for carry forward to offset future taxable income. Year Ended Year Ended December 31, 2018 December 31, 2019 United States federal statutory rate 21.00% 21.00% Effective income tax rate 1.99% -0.01% Question: What was the 2019 percentage change in pre-tax losses? Response #: Domain Knowledge: Pre-tax losses, or operating losses, refer to financial losses that a company incurs before considering the effects of income taxes. To find the 2019 percentage change in pre-tax losses, we need to find the difference between the new and the old value of the pre-tax losses, then divide the obtained difference by the old value and multiply this value by 100. Gold Evidences: The pre-tax losses in 2019 are \$19,573.
The pre-tax losses in 2018 are \$25,403.
Steps: 1. Subtract 25403 from 19573. Response 1: 19573 - 25403 = -5830 2. Divide #1 by 25403. Response 2: $\frac{-5830}{25403}=-0.2295$ 3. Multiply #2 by 100. Response 3: $0.2295\times 100=-22.95$ The final answer is -22.95%.
@@
Basel III Standardized Transitional Basel III Advanced Transitional (in millions) Dec 31, 2017 Dec 31, 2016 Dec 31, 2017 Dec 31, 2016 Regulatory capital CET1 capital \$184,375 \$179,319 \$184,375 \$179,319 Tier 1 capital(a) \$184,375 \$179,341 \$184,375 \$179,341 Total capital \$195,839 \$191,662 \$189,419 \$184,637 Assets Capital ratios(c) CET1 13.8% 13.7% (e) 15.0% Tier 1(a) 13.8% 13.7% (e) 15.0% Total 14.7% 14.6% (e) 15.4% Tier 1 leverage(d) 8.7% 8.6% 8.7%
Question: What is the sum of CET1 capital, Tier 1 capital, and Total capital in 2017? (in million)
Response #:
Domain Knowledge: In banking, CET1 capital, Tier 1 capital, and Total capital are used to measure a bank’s financial health and stability. CET1 (Common Equity Tier 1) capital consists mainly of common shares and retained earnings. Tier 1 capital is the core capital mainly comprising shareholders’ equity and disclosed reserves. Total capital includes both Tier 1 capital and Tier 2 capital, which consists of hybrid capital instruments and subordinated debt.
Gold Evidences:
Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional CET1 capital on Dec 31, 2017 as \$184,375 million. Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional Tier 1 capital on Dec 31, 2017 as \$184,375 million. Table shows JPMorgan Chase Bank, N.A. Basel III Standardized Transitional Total capital on Dec 31, 2017 as \$195,839 million.
Steps:
1. Add CET1 capital and Tier 1 capital. Response 1: \$184,375 + \$184,375 = \$368,750 2. Add #1 and Total capital. Response 2: #1 + \$195,839 = \$564,589
The final answer is \$564,589 million.
Ground-Truth Steps: \$184,375 + \$184,375 + \$195,839 + \$184,375 + \$184,375 + \$189,419 = \$1,122,758
Reason: The model missed adding JPMorgan Chase Bank, N.A. Basel III Advanced Transitional CET1 capital (\$184,375), Tier 1 (\$184,375), and Total Capital (\$189,419) in 2017.
@@
2013 2014 2015 2016 2017 Masco \$138.48 \$155.26 \$200.79 \$227.08 \$318.46 S&P 500 Index \$132.04 \$149.89 \$151.94 \$169.82 \$206.49 S&P Industrials Index \$140.18 \$153.73 \$149.83 \$177.65 \$214.55 S&P Consumer Durables & Apparel Index \$135.84 \$148.31 \$147.23 \$138.82 \$164.39 Question: What was the difference in percentage cumulative total shareholder return on Masco common stock versus the S&P 500 Index for the five-year period ended 2017? Response #: Domain Knowledge: Cumulative total shareholder return is a measure of the overall performance of a stock over a given period of time. It includes the change in the stock’s price as well as any dividends paid to shareholders during that time. Gold Evidences: The cumulative total shareholder return for Masco in 2017 is \$318.46. The cumulative total shareholder return for the S&P 500 Index in 2017 is \$206.49. Steps: 1. Divide the cumulative total shareholder return for Masco in 2017 (318.46) by the cumulative total shareholder return for the S&P 500 index in 2017 (206.49) Response 1: 318.46 / 206.49 = 1.5422 2. Subtract 1 from the quotient obtained in #1 Response 2: 1.5422 - 1 = 0.5422 3. Multiply the value obtained in #2 with 100 to find the percentage difference Response 3: 0.5422 * 100 = 54.22 The final answer is 54.22%. Ground-truth steps: [((318.46 - 100.0)/100.0)*100.0 - ((206.49 - 100.0)/100.0)*100.0] = 111.97 Reason: Wrong formula for return on investment